The Gentry Garden center sells 10,000 bags of lawn fertilizer annually. Each bag costs the firm $2.00, inventory carrying cost is 20%, and the cost of placing an order with the supplier is $20. Answer the following questions: 1) What is the firm’s Economic Ordering Quantity? 2) What is the firm’s average inventory? 3) How many times must the firm order inventory in a year?
Answer #1

1

Economic order quantity = ((2*annual demand quantity* fixed cost per order)/ annual holding cost per unit)^(0.5)

=((2*10000*20/(2*0.2)))^0.5

=1000

2

Average inventory level = EOQ/2 = 1000/2 = 500

3

Number of orders = sales/EOQ = 10000/1000=10